OpenStudy (anonymous):

Find the vertex and axis of symmetry of the graph of the function ..... F(x) = x^2 + 4x + 3

OpenStudy (campbell_st):

the axis of symmetry is found for for a parabola in the general form \[y = ax^2 + bx + c\] by using the formula \[x = \frac{-b}{2a}\] the vertex will be \[(\frac{-b}{2a}, f(\frac{-b}{2a}))\] in your question a = 1 and b = 4

OpenStudy (anonymous):

thank you!